# Analysis of Regular Sequences: Summatory Functions and Divide-and-Conquer Recurrences

### v1

@@ 1.1 Overview
Studying the summatory function was motivated by the fact that in several well-known examples of regular sequences, the sequences themselves are fluctuating too much so that it is impossible to establish a good asymptotic expansion for the regular sequence itself. For instance, for the binary sum of digits function $s_{2}$, we have $s_{2}(2^{k}-1)=k$ and $s_{2}(2^{k})=1$ for all integers $k\geq 0$, so the most precise asymptotic expansion for $s_{2}(n)$ is $s(n)=O(\log n)$ for $n\to\infty$. However, the summatory function might admit a good asymptotic expansion: For the summatory function of the binary sum of digits function, we have $\sum_{0\leq n<N}s_{2}(n)=\frac{1}{2}N\log_{2}N+N\Phi(\log_{2}N)$ for some 1-periodic continuous function $\Phi$ as $N\to\infty$; see Delange [3]. In this particular example, there is no error term; in general, an error term is to be expected.
@@
For other regular sequences, the sequence itself might admit a good asymptotic expansion. One example are sequences associated with divide-and-conquer schemes [11, 10], for example, the worst case analysis of the number of comparisons in the merge sort algorithm. These are closely related to the so-called “master theorems”; see the discussion in [10]. These sequences are easily seen [10, Equation (2.1)] to be regular sequences (as long as the toll function is regular). While Hwang, Janson, and Tsai [10] provide a direct proof for the asymptotic behaviour and give plenty of examples, the question is whether these results can also be obtained by using the results in [7]. In the present paper, we see such a sequence as the summatory function of its forward difference, and we show that for polynomial toll functions, we get a good asymptotic expansion in the vast majority of cases. The result is formulated in Theorems 1.7 and 1.8. In contrast to [10], we are not constrained to cases where the toll function is asymptotically smaller than the sequence and Fourier coefficients can be computed using the results of [7].
The remaining paper is structured as follows. In Section 1.2, we recall the definition and the relevant results on regular sequences. This is followed in Section 1.3 by the statement of our new result on the $k$-fold summatory function. In Section 1.4, we present the state of the art for divide-and-conquer sequences and state our version the result in Section 1.5. Sections 2 and 3 are devoted to the proofs of our theorems. We conclude with an explicit example in Section 4.
@@ 1.5 Divide-and-Conquer Sequences: Polynomial Toll Function
It is not hard to see that for divide-and-conquer sequences $g$, a sequence satisfying the recurrence (5) can be seen as a regular sequence: We have
@@
This theorem is also proved in Appendix A.
@@
3 Divide-and-Conquer Recurrences: Proof of Theorem 1.7
@@
$x(2n+1)$ $=\alpha x(n)+\beta x(n+1)+g(2n+1)+((1-\beta)x(1)-g(1))\delta_{0}(n).$
@@
$\mu_{r}$$\coloneqq\begin{cases}\beta&\text{ if }r=0,\\ \alpha&\text{ if }r=1,\end{cases}$ $A_{r}$ $\coloneqq\begin{pmatrix}\mu_{r}&b_{r}&d_{r}\\ 0&\widetilde{A}_{r}&0\\ 0&0&\llbracket r=0\rrbracket\end{pmatrix},$
@@
4 Example
We conclude with one example to illustrate the results.
Example 4.1.
Consider the divide-and-conquer algorithm for finding the minimum and the maximum of a list of $n$ elements. The number $x(n)$ of comparisons needed satisfies (5) for $n\geq 3$ with $a=b=1$ and $g(n)=2$ for $n\geq 3$ and with $x(1)=0$ and $x(2)=1$; cf. [11, Example 3.2].
By Theorem 1.8 (and [8, Theorem B] to deal with the fact that the divide-and-conquer recurrence is only valid for $n\geq 3$ instead of $n\geq 2$; see Appendix B for details), we get
$x(n)=n\Phi(\{\log_{2}n\})+O(n^{\varepsilon})$
Appendix A Proof of Theorem 1.8
Proof A.1 (Proof of Theorem 1.8).
@@ 4 Example
We can now use the SageMath code [^4] The code is available at https://gitlab.com/dakrenn/regular-sequence-fluctuations.
accompanying [7].

### v2

@@ 1.1 Overview
Studying the summatory function was motivated by the fact that in several well-known examples of regular sequences, the sequences themselves are fluctuating too much so that it is impossible to establish a good asymptotic expansion for the regular sequence itself. For instance, for the binary sum of digits function $s_{2}$, we have $s_{2}(2^{k}-1)=k$ and $s_{2}(2^{k})=1$ for all integers $k\geq 0$, so the most precise asymptotic expansion for $s_{2}(n)$ is $s_{2}(n)=O(\log n)$ for $n\to\infty$. However, the summatory function might admit a good asymptotic expansion: For the summatory function of the binary sum of digits function, we have $\sum_{0\leq n<N}s_{2}(n)=\frac{1}{2}N\log_{2}N+N\Phi(\log_{2}N)$ for some 1-periodic continuous function $\Phi$ as $N\to\infty$; see Delange [3]. In this particular example, there is no error term; in general, an error term is to be expected.
@@
For other regular sequences, the sequence itself might admit a good asymptotic expansion. One example are sequences associated with divide-and-conquer schemes [10, 11], for example, the worst case analysis of the number of comparisons in the merge sort algorithm. These are closely related to the so-called “master theorems”; see the discussion in [11]. These sequences are easily seen [11, Equation (2.1)] to be regular sequences (as long as the toll function is regular). While Hwang, Janson, and Tsai [11] provide a direct proof for the asymptotic behaviour and give plenty of examples, the question is whether these results can also be obtained by using the results in [7]. In the present paper, we see such a sequence as the summatory function of its forward difference, and we show that for polynomial toll functions, we get a good asymptotic expansion in the vast majority of cases. The result is formulated in Theorems 1.7 and 1.8. In contrast to [11], we are not constrained to cases where the toll function is asymptotically smaller than the sequence and Fourier coefficients can be computed using the results of [7].
The remaining paper is structured as follows. In Section 1.2, we recall the definition and the relevant results on regular sequences. This is followed in Section 1.3 by the statement of our new result on the $k$-fold summatory function. In Section 1.4, we present the state of the art for divide-and-conquer sequences and state our version the result in Section 1.5. An explicit example is discussed in Section 1.6. Sections 2 and 3 are devoted to the proofs of our theorems.
@@ 1.5 Divide-and-Conquer Sequences: Polynomial Toll Function
For divide-and-conquer sequences $g$, a sequence satisfying the recurrence (5) can be seen as a regular sequence: It is not hard to see that we have
@@
This theorem is also proved in Section 3.
1.6 Example
We conclude this introductory section with one example to illustrate the results.
Example 1.9.
Consider the divide-and-conquer algorithm for finding the minimum and the maximum of a list of $n$ elements. The number $x(n)$ of comparisons needed satisfies (5) for $n\geq 3$ with $a=b=1$ and $g(n)=2$ for $n\geq 3$ and with $x(1)=0$ and $x(2)=1$; cf. [10, Example 3.2].
By Theorem 1.8 (and [8, Theorem B] to deal with the fact that the divide-and-conquer recurrence is only valid for $n\geq 3$ instead of $n\geq 2$; see Section 4 for details), we get
$x(n)=n\Phi(\{\log_{2}n\})+O(n^{\varepsilon})$
@@ 1.6 Example
3 Divide-and-Conquer Recurrences: Proof of Theorems 1.7 and 1.8
@@
$x(2n+1)$ $=\alpha x(n)+\beta x(n+1)+g(2n+1)+((1-\beta)x(1)-g(1))\delta_{0}(n)$
@@
$\mu_{r}$$\coloneqq\begin{cases}\beta&\text{ if }r=0,\\ \alpha&\text{ if }r=1,\end{cases}$ $A_{r}$ $\coloneqq\begin{pmatrix}\mu_{r}&b_{r}&d_{r}\\ 0&\widetilde{A}_{r}&0\\ 0&0&\llbracket r=0\rrbracket\end{pmatrix}$
@@
Proof 3.9 (Proof of Theorem 1.8).
@@
We can now use SageMath [^4] The code for this example is available at https://arxiv.org/src/2403.06589/anc; it uses the code accompanying [7] which is available at at https://gitlab.com/dakrenn/regular-sequence-fluctuations. to compute Fourier coefficients and to produce Figure 1.